# Review of nonflow estimation methods and uncertainties in relativistic heavy-ion collisions

### v2

@@ 2.2.1 Inverse multiplicity scaling method
    it assumes the $1/N$ scaling of nonflow correlations, implying no change in the physics of nonflow correlations from low- to high-multiplicity collisions, a strict proportionality to $N$ of the abundance of nonflow sources, and no selection biases for those low-multiplicity events;
@@
The strong assumption of ii) and the fact that how nonflow correlations vary with the collision centrality/multiplicity make it difficult to assess the robustness and the associated uncertainties of the estimated nonflow by the simple $1/N$ scaling method.
2.2.2 Near-side jet yield scaled subtraction
Nuclear effects, like jet quenching, result in modifications to jet-correlations [24], thus nonflow effects change from peripheral to central collisions. Particle production mechanisms and distributions change, for example, yielding the baryon-over-meson enhancement from peripheral to central collisions [54, 55, 56], which would modify the nonflow effects from resonance decays. Any those changes will cause nonflow to deviate from the simple $1/N$ scaling.
Both jet-correlations and resonance decays yield a near-side correlation peak at $(\Delta\eta,\Delta\phi)\sim(0,0)$. Dijets contribute to an away-side correlation at $\Delta\phi\sim\pi$ but more or less uniform in $\Delta\eta$. One may take the difference between small $\Delta\eta$ (short-range) and large $\Delta\eta$ (long-range) correlations, properly normalized, to arrive at a near-side correlated yield (primarily composed of contributions from resonance decays and intra-jet fragments). One can then compare this near-side yield in high-multiplicity events to that in low-multiplicity events. Any difference would indicate modification of nonflow correlations from low- to high-multiplicity collisions, multiplicity selection biases [47, 48], or likely both. One may take the ratio of the correlated near-side yields ($Y_{\rm high}/Y_{\rm low}$) as a scaling factor to apply on to the low-multiplicity nonflow in Eq. (7) to possibly take into account modifications to nonflow correlations [57, 48, 58]. The estimated flow in high-multiplicity events would then be
$V_{2}=V_{2}^{\rm high}-\frac{N_{\rm low}}{N_{\rm high}}\frac{Y_{\rm high}}{Y_{\rm low}}V_{2}^{\rm low}\,.$ (8)
Such a scaling, while an improvement to the simple $1/N$ scaling, comes with its own issues:
@@ 2.2.2 Near-side jet yield scaled subtraction
    the near-side nonflow correlation shape may change/broaden, the effect of which is not included by the simple scaling of the correlated yields;
@@
    the away-side jet-correlated yield unlikely scales with the near-side one because of “trigger” biases (for example, surface bias at high $p_{T}$), and even if it scales, the scaling factor $Y_{\rm high}/Y_{\rm low}$ includes resonance decays besides near-side jet contribution and is therefore already an incorrect scaling factor, needless to say that the away-side correlation shape can be significantly modified because of large jet-quenching effects;
@@
    the correlated yield analysis is difficult at low $p_{T}$ because of large combinatorial background.
@@ 2.2.3 Template fit
In the template fit method [59, 18], the two-particle correlation is assumed to be composed of a scaled correlation from low-multiplicity events (considered all as nonflow) and a series of Fourier harmonics (to represent flow) except the $n=1$ component. This is illustrated in Fig. 3. Namely, the following function is fit to two-particle correlations in high-multiplicity events,
@@
where $F$ is a fit parameter to scale the low-multiplicity correlations. Note that in the harmonic series, the first harmonic is excluded. In other words, the fitting constraint is to assume that the first harmonic is all nonflow; there is no flow harmonic of the first order. (Of course, without any imposed constraint, the fit would not yield any unique result because anything can be described by a Fourier series.) In effect, the template fit method assumes the nonflow component in high-multiplicity events to scale according to $V_{1}$. If $V_{1}$ is inversely proportional to multiplicity (which is approximately the case, see e.g. Ref. [58]), then the template fit method is similar to the $1/N$ scaling method in Sect. 2.2.1. Any deviation of the fit parameter $F$ from the $1/N$ scaling factor can be considered as a modification in nonflow correlations from low to high multiplicity, taken care of by a constant scaling without change in shape. This is similar to the near-side jet yield scaled subtraction described in Sect. 2.2.2, except that the scaling here is determined by the $V_{1}$ dipole component (mainly the away-side correlation, probably dominated by global momentum conservation [28]) instead of the near-side correlation.
Figure 3: Sketch of the template fit method and the dipole method. The heavy-ion pair correlations are considered to be composed of nonflow correlations scaled from pp and flow correlations. In the template fit, the extracted flow magnitudes are with respect to the pair multiplicity between the two curves, i.e., the nonflow pairs scaled from pp is first subtracted. In the dipole method, the flow magnitudes are with respect to the total pair multiplicity underneath the upper curve. The methods also apply to small system analysis, where the label “heavy-ion” is to be replaced by “high-multiplicity events” and “scaled pp” by “low-multiplicity events.”
@@
Eq. (13) is similar to Eqs. (7,8) except the denominator. The denominator comes about because the normalization $G$ in Eq. (12) is not all the pairs in the high-multiplicity event but that excluding those nonflow particle pairs. In other words, the physics picture is that the high-multiplicity event is composed of nonflow particles (the white area in Fig. 3) and flow particles (the hatched area in Fig. 3) and the nonflow particles do not participate in the flow.
The template fit method attempts to account for nonflow changes by scaling the correlations in low-multiplicity events. It is similar in spirit to the near-side jet yield scaled subtraction method, differing in the assumption of how nonflow changes. So a similar set of assumptions and shortcomings, namely:
@@
  • iv)
    it assumes that the nonflow pairs do not participate in the final-state collective flow, and the extracted flow is relative to the reduced average pair multiplicity–the normalization $G$ in Eq. (9) or the area between the two curves in Fig. 3–after subtraction of those nonflow pairs. However, the cross pairs (one from a nonflow source and the other from the collective flow bulk) are included in the normalization $G$ and thus are considered as part of flow. This normalization issue is mostly a small effect but may be relevant for small-system flow studies.
The last assumption is unique to the template fit method. For all other nonflow estimation methods, the final extracted flows are based on the total event (pair) multiplicity, including nonflow particles–see bullet iii) in Sect. 2.2.1 (not explicitly stated for other nonflow estimation methods).
2.2.4 Dipole method
This method [18, 19] is very similar to the template fit method in Sect. 2.2.3, also assuming the first harmonic $V_{1}$ to be all nonflow, and the $V_{1}$ in low-multiplicity events is scaled to match that in high-multiplicity events, as in Eq. (11). The $V_{n}^{\rm sub}$ is, however, defined in terms of the total pair multiplicity $N_{\rm pair}^{\rm high}$, not $G$ as in Eq. (12). Namely, $N_{\rm pair}^{\rm high}V_{n}^{\rm high}=FN_{\rm pair}^{\rm low}V_{n}^{\rm low}+N_{\rm pair}^{\rm high}V_{n}^{\rm sub}$, which leads to
@@ 2.2.4 Dipole method
This assumes effectively that all those nonflow particles, while keeping their inter-particle nonflow correlations, have gained collective flow themselves individually. This is as same as the assumption made in Sects. 2.2.1 and 2.2.2. In the template fit method, on the other hand, the low-multiplicity correlated yield is scaled and subtracted first, and those particles are not counted as part of collective flow as aforementioned. Effectively, the dipole method is as same as the template method except that the scaled low-multiplicity correlation is first lowered to average zero. The dipole method is sometimes called “$c_{1}$ method” (where $c_{1}\equiv V_{1}$[18, 19].
@@ 2.3 Data-driven fitting methods
As seen from Sects. 2.1 and 2.2, nonflow contamination is difficult to remove, involving strong assumptions. Another approach to estimate nonflow contamination is quite different from those described in Sects. 2.1 and 2.2, and is data driven. It relies on correlation structures observed in data, and identifies local peaks and ridges to attribute to nonflow correlations and the more or less $\Delta\eta$-independent underlying distribution to collective flow. It then involves fitting the correlation data with pre-defined functional forms in a data-driven way.
The advantage of these fitting methods is that they are data driven, having minimal assumptions about the physics of nonflow and no reliance on the evolution of nonflow over centrality/multiplicity. The downside, however, is that the fit functional forms to nonflow correlations are ad hoc and that the fitting can sometimes be tedious.
@@
One data-driven technique is to perform 2-dimensional $(\Delta\eta,\Delta\phi)$ fits to two-particle correlations [60, 44, 61, 62, 63]. Raw two-particle correlations are predominantly of a triangular shape in $\Delta\eta$, which comes from the approximately uniform single-particle density distribution within limited $\eta$ acceptance (for example, $|\eta|<1$ in the STAR experiment [64]). The triangular acceptance can be largely corrected by the mixed-event technique, resulting in two-particle $(\Delta\eta,\Delta\phi)$ correlations, such as the one sketched in Fig. 1, that are dominated by an overall pedestal. The fine structures over the pedestal are nonflow correlations and a flow modulation along $\Delta\phi$ (with possibly a $\Delta\eta$ dependence). These features have been generally described in the introduction.
The near-side nonflow peaks can be modeled by 2D Gaussians, and the flow modulation is described by a Fourier series. The Fourier coefficients, corresponding to flow harmonics, are typically assumed to be $\Delta\eta$ independent. The two-particle correlations can then be fit with a functional form including Gaussians and Fourier harmonics. This 2D fitting method is data-driven; the nonflow correlation shapes are dictated by the data structure and modeled. The method fits the data in a given centrality or multiplicity bin and does not rely on assumptions using low-multiplicity baseline events. A recent 2D fit study indicates an approximately 40% nonflow in central isobar collisions [62, 63], in line with the central Au+Au data considering multiplicity dilution of nonflow.
It is noteworthy that the Fourier flow harmonics can be $\Delta\eta$ dependent, for example, due to flow decorrelations [34, 35, 65, 36]. In addition, the $v_{n}$, although long ranged, can be $\eta$ dependent [32]; such a dependence can result in $\Delta\eta$-dependent Fourier coefficients as well. Including these possible $\eta$-dependencies of the flow harmonics as free parameters in the fit model would not be fruitful as the correlation data would not have sufficient constraining power over these parameters (including those modeling nonflow). Usually the effects due to the possible $\eta$-dependent flow harmonics are assessed as part of the systematic uncertainties.
However, the information of the $\Delta\eta$-dependent harmonics can be obtained from other analyses and can be factored into the 2D fit function. We leave discussions on this to the next section.
@@
The spirit of the 1D fitting method is similar to that of the 2D fitting–decomposing the flow and nonflow in the inclusive measurement of $V_{n}(\Delta\eta)$ as a function of $\Delta\eta$. Figure 4 illustrates a typical $V_{2}(\Delta\eta)$ dependence on $\Delta\eta$. It is generally a decreasing function of $\Delta\eta$, primarily because of the short-range nonflow contributions. For tracking detectors, track merging is sometimes important, which creates a dip at $\Delta\eta\sim 0$ as indicated by the dashed curve in Fig. 4. When track merging effects are significant, the average $V_{n}$ from full-event analysis can be even smaller than subevent analysis which automatically excludes the dip. In such a case, the difference between full-event and subevent $V_{n}$ in the $\Delta\eta$-gap methods would not be a correct assessment of systematic uncertainties.
Compared to the 2D fitting method in Sect. 2.3.1, 1D fitting is more straightforward because the $\Delta\phi$ dimension is collapsed in the $V_{n}$ measurement. It is also easier to include the possible $\Delta\eta$ dependence of flow. There are two sources for $\Delta\eta$ dependence of flow as aforementioned: $\eta$-dependent single-particle flow and flow $\Delta\eta$ decorrelations.
The single particle flow $v_{n}$ can depend on $\eta$ [32]. Such a dependence can be examined, for the case of $n=2$ for example, by four-particle cumulant measurement $v_{2}\{4\}(\eta)$ where nonflow is largely eliminated, and by $v_{2}\{{\textsc{zdc}}\}(\eta)$ with respect to the first order harmonic plane of spectator neutrons measured by Zero-Degree Calorimeters (ZDC) which is free of nonflow. Note that $v_{2}\{{\textsc{zdc}}\}$ measures approximately the average $\langle{v_{2}}\rangle$ and $v_{2}^{2}\{4\}\approx\langle{v_{2}}\rangle^{2}-\sigma^{2}$ where $\sigma$ denotes Gaussian width of $v_{2}$ fluctuations, whereas the flow component in the two-particle measurement $V_{2}$ is $v_{2}^{2}\{2\}=\langle{v_{2}}\rangle^{2}+\sigma^{2}$ [16]. As an approximation, one may obtain $v_{2}\{2\}(\eta)$ from $v_{2}\{4\}(\eta)$ and $v_{2}\{{\textsc{zdc}}\}(\eta)$. Alternatively, one may make the reasonable assumption that the effect of $v_{2}$ fluctuations is proportional to $\langle{v_{2}}\rangle$ and apply a scale factor as a free fit parameter to $v_{2}\{4\}(\eta)$ or $v_{2}\{{\textsc{zdc}}\}(\eta)$ to obtain $v_{2}\{2\}(\eta)$. The $\Delta\eta$ dependence can be obtained by folding two single-particle distributions of $v_{2}(\eta_{1})$ and $v_{2}(\eta_{2})$.
Another source of $\Delta\eta$ dependence is the flow decorrelation, $r_{n}(\Delta\eta)\approx 1-F_{n}\Delta\eta$ [65, 36]. Such decorrelations can result from fluctuations of the harmonic flow magnitude as well as harmonic planes over $\Delta\eta$ [66]. The decorrelation parameter $F_{n}$ of a few percent has been measured at RHIC [67] and the LHC [65, 36].
The $\Delta\eta$ dependence of the flow component in two-particle $V_{2}(\Delta\eta)$ is the combination of the two effects above. This is illustrated by the magenta curve in Fig. 4. The nonflow component in $V_{2}(\Delta\eta)$ is modeled by Gaussians in $\Delta\eta$. Fits are applied to $V_{2}(\Delta\eta)$ excluding the track merging $\Delta\eta\sim 0$ region, or alternatively including a negative Gaussian at $\Delta\eta=0$ to model it, treating the flow fluctuation scale factor as one parameter in the fit. The blue curve is an illustration of the fitted flow, scaled up by the fit parameter from the shape given in magenta. The idea of the fit is to have the best description of $V_{2}(\Delta\eta)$ with the known shape of flow and Gaussian-shaped nonflow.
@@
Figure 4: Illustration of the 1D fit method. The black curve indicates a $V_{2}$ measurement, where the dip at $\Delta\eta\sim 0$ indicated by the dashed curve is caused by track merging in a typical tracking detector. The magenta curve incorporates $\Delta\eta$ dependencies from single-particle $v_{2}(\eta)$ and flow decorrelation $1-F_{n}\Delta\eta$. The blue curve indicates the flow component in $V_{2}$ scaled up from the magenta curve to include flow fluctuations, with the scaling factor treated as a fit parameter. The difference between the black curve and the blue curve is nonflow contribution to the $V_{2}$ measurement. The vertical axis is zero-suppressed and the values are only order-of-magnitude indications.

### v4

@@ 2.2.1 Inverse multiplicity scaling method
    it assumes the $1/N$ scaling of nonflow correlations, implying no change in the physics of nonflow correlations from low- to high-multiplicity collisions, a strict proportionality to $N$ of the abundance of nonflow sources, and no selection biases for those low- (and high-)multiplicity events;
@@
The strong assumption of ii) and the fact that it is a priori unknown how nonflow correlations vary with the collision centrality/multiplicity make it difficult to assess the robustness and the associated uncertainties of the estimated nonflow by the simple $1/N$ scaling method.
2.2.2 Near-side jetlike yield-scaled subtraction
As noted in the Sect. 2.1.1, nuclear effects, like jet quenching, result in modifications to jet correlations [43, 44], thus nonflow effects change from peripheral to central collisions. Likewise, particle production mechanisms and distributions change with centrality, yielding, for example, the baryon-over-meson enhancement from peripheral to central collisions [115, 116, 117], which would modify the nonflow effects from resonance decays. Any of those changes will cause nonflow to deviate from the simple $1/N$ scaling.
Both jetlike correlations and resonance decays produce a near-side correlation peak at $(\Delta\eta,\Delta\phi)\sim(0,0)$. Dijets contribute to an away-side correlation at $\Delta\phi\sim\pi$ but more or less uniform in $\Delta\eta$. One may take the difference between small-$\Delta\eta$ (short-range) and large-$\Delta\eta$ (long-range) correlations, properly normalized, to arrive at a near-side correlated yield, which is primarily composed of contributions from resonance decays and intra-jet fragments. One can then compare this near-side yield in high-multiplicity events to that in low-multiplicity events. Any difference would indicate modification of nonflow correlations from low- to high-multiplicity collisions and/or multiplicity selection biases [62, 64, 112, 113, 114], or likely both. One may take the ratio of the correlated near-side yields ($Y_{\rm high}/Y_{\rm low}$) as a scaling factor to apply onto the low-multiplicity nonflow $V_{n}^{\rm low}$ in Eq. (7) to possibly take into account modifications to nonflow correlations [61, 113, 68, 118]. The estimated flow in high-multiplicity events would then be
$V_{n}^{\rm sub}=V_{n}^{\rm high}-\frac{Y_{\rm high}}{Y_{\rm low}}\frac{N_{\rm low}}{N_{\rm high}}V_{n}^{\rm low}\,.$ (8)
Such a scaling, while an improvement over the simple $1/N$ scaling, comes with its own issues:
@@ 2.2.2 Near-side jetlike yield-scaled subtraction
    the near-side nonflow correlation shape may change/broaden, the effect of which is not included by the simple scaling of the correlated yield;
@@
    the away-side jet correlated yield unlikely scales with the near-side one because of “trigger” biases (for example, surface bias at high $p_{T}$), and even if it scales, the scaling factor $Y_{\rm high}/Y_{\rm low}$ includes resonance decays besides near-side jet contribution and is therefore already an incorrect scaling factor, needless to say that the away-side correlation shape can be significantly modified because of large jet-quenching effects;
@@
    the correlated yield analysis may be difficult at low $p_{T}$ because of the subtraction of large combinatorial background.
The near-side yield-scaled subtraction is usually done at the Fourier coefficient level by Eq. (8). Effectively, it is a subtraction of the scaled correlation function in low-multiplicity events from the correlation function in high-multiplicity events, ignoring any changes in the correlation shape. Shape changes in nonflow correlations noted in point i) above can, to some extent, be inspected by the small- and large-$\Delta\eta$ correlation difference and can in principle be considered in nonflow subtraction. Of course, this difference would be purely nonflow correlations only under the assumption that flow is $\Delta\eta$ independent. If flow is $\Delta\eta$ dependent, then it becomes difficult to disentangle the two effects. Moreover, the approximate $\Delta\eta$-independent away-side correlations are also subtracted in such a difference, so no information on any shape change in $\Delta\phi$ correlations is retained. The ultimate approach is to use data-driven fitting methods, which will be discussed in Sect. 2.3.
Two-particle cumulant flow involves particle pairs. The nonflow correlation shapes are often studied in terms of per-trigger normalized correlated yield. Besides the nonflow correlation shape, the number of trigger particles also matters for flow. If the number of trigger particles does not scale with event multiplicity, then the $1/N$ scaling will be invalid and this is certainly the case for jet correlations. For most low-$p_{T}$ flow analysis in heavy-ion collisions, this is not an issue because the trigger particle is the charged hadron itself. However, this is an issue for small system collisions like $pp$ because of multiplicity selection biases. The number of minijets is biased and will not be proportional to the event multiplicity. This imposes a large uncertainty on small-system flow analysis.
@@ 2.2.3 Template fit
In the template fit method [119, 65, 31, 32], the two-particle correlation is assumed to be composed of a scaled correlation from low-multiplicity events (considered all as nonflow) and a series of Fourier harmonics (to represent flow) except the $n=1$ component. This is illustrated in Fig. 3. Namely, the following function is fit to two-particle correlations in high-multiplicity events,
@@
where $F$ is a fit parameter to scale the low-multiplicity correlations (template). Note that in the harmonic series, the first harmonic is excluded. In other words, the fitting constraint is to assume that the first harmonic is all nonflow; there is no flow harmonic of the first order. [^3] Of course, without any imposed constraint, the fit would not yield any unique result because anything can be described by a Fourier series.
The template fit assumes, effectively, that the nonflow component in high-multiplicity events scales according to $V_{1}$. If $V_{1}$ is inversely proportional to multiplicity (which is approximately the case, see e.g. Ref. [68]), then the template fit method is similar to the $1/N$ scaling method in Sect. 2.2.1. In other words, the assumption here is that nonflow scales with multiplicity in the same way as $V_{1}$. Any deviation of the fit parameter $F$ from the $1/N$ scaling factor can be considered as a modification in nonflow correlations from low to high multiplicity, taken care of by a constant scaling without change in shape. This is similar to the near-side jetlike yield-scaled subtraction described in Sect. 2.2.2, except that the scaling here is determined by the $V_{1}$ dipole component (mainly the away-side correlation, probably dominated by global momentum conservation [47]) instead of the near-side correlation. The template fit can therefore be coined as a $V_{1}$-scaled subtraction method, just like the near-side scaled subtraction can be recasted as a fitting method to match the near-side yield from small- and large-$\Delta\eta$ difference.
Figure 3: Sketch of the template fit method and the dipole scaling method. The heavy-ion pair correlations are considered to be composed of nonflow correlations scaled from $pp$ and flow correlations. In the template fit, the extracted flow magnitudes are with respect to the pair multiplicity between the two curves, i.e., the nonflow pairs scaled from $pp$ is first subtracted. In the dipole scaling method, the flow magnitudes are with respect to the total pair multiplicity underneath the upper curve. The methods also apply to small system analysis, where the label “heavy-ion” is to be replaced by “high-multiplicity events” and “scaled $pp$” by “low-multiplicity events.”
@@
Eq. (13) is similar to Eqs. (7, 8) except the denominator. The denominator comes about because the normalization $G$ in Eq. (12) does not include all the pairs in the high-multiplicity events but excludes those genuine nonflow particle pairs. It assumes that the nonflow pairs do not participate in the final-state collective flow, and the extracted flow is relative to the reduced average pair multiplicity (the underlying event) excluding those nonflow pairs. The physics picture is that the high-multiplicity event is composed of nonflow particles (the white area underneath the lower curve in Fig. 3) and flow particles (the area between the two curves in Fig. 3). However, the cross pairs (one from a nonflow source and the other from the collective flow bulk) are included in the normalization $G$ and thus are considered as part of flow. This underlying event normalization issue is mostly a small effect but may be relevant for small-system flow studies. To some extent this is an intrinsic uncertainty as it is not a priori clear which picture is more realistic. This normalization difference is unique to the template fit method. For all other nonflow estimation methods, the final extracted flows are based on the total event (pair) multiplicity, including nonflow particles–see bullet iii) in Sect. 2.2.1 (not explicitly stated for other nonflow estimation methods).
The template fit method attempts to account for nonflow changes by scaling the correlations in low-multiplicity events. It is similar in spirit to the near-side jetlike yield-scaled subtraction method, differing in the assumption of how nonflow changes. So a similar set of assumptions and shortcomings, namely:
@@
2.2.4 Dipole scaling method
The dipole scaling method [31, 32] is very similar to the template fit method in Sect. 2.2.3, also assuming the first harmonic $V_{1}$ to be all nonflow, and the $V_{1}$ in low-multiplicity events is scaled to match that in high-multiplicity events, as in Eq. (11). The $V_{n}^{\rm sub}$ is, however, defined in terms of the total pair multiplicity $N_{\rm pair}^{\rm high}$, not $G$ as in Eq. (12). Namely, $N_{\rm pair}^{\rm high}V_{n}^{\rm high}=FN_{\rm pair}^{\rm low}V_{n}^{\rm low}+N_{\rm pair}^{\rm high}V_{n}^{\rm sub}$, which leads to
@@ 2.2.4 Dipole scaling method
This assumes effectively that all those nonflow particles, while keeping their inter-particle nonflow correlations, have attained collective flow themselves individually. This is as same as the assumption made in Sect. 2.2.1 and 2.2.2. In the template fit method, on the other hand, the low-multiplicity correlated yield is scaled and subtracted first, and those particles are not counted as part of collective flow as aforementioned. Effectively, the dipole method is as same as the template method except that the subtracted average baseline in the low-multiplicity template is added back to the underlying event. The dipole method is sometimes called “$c_{1}$ method” (where $c_{1}\equiv V_{1}$[31, 32].
@@ 2.3 Data-driven fitting methods
As seen from Sect. 2.1 and 2.2, nonflow contamination is difficult to remove, involving strong assumptions. Another approach to estimate nonflow contributions is quite different from those described in Sect. 2.1 and 2.2. It relies on correlation structures observed in data, and identifies local peaks and ridges to attribute to nonflow correlations and the approximately $\Delta\eta$-independent underlying distribution to collective flow. It then involves fitting the correlation data with pre-defined functional forms in a data-driven way.
The advantage of these fitting methods is that they are data driven, having minimal assumptions about the physics of nonflow and no reliance on the evolution of nonflow over centrality/multiplicity. The downside, however, is that the fit functional forms for nonflow correlations are ad hoc and one needs to ensure that the nonflow functional forms are reasonable descriptions of the data by trial and error, examining fit qualities by using, e.g. the fit $\chi^{2}/{\rm ndf}$ values.
@@
One data-driven approach is to perform two-dimensional (2D) fits to two-particle $(\Delta\eta,\Delta\phi)$ correlations [120, 105, 121, 122, 123]. Raw two-particle correlations are predominantly of a “triangle” shape in $\Delta\eta$, which comes from the approximately uniform single-particle density distribution within limited $\eta$ acceptance (for example, $|\eta|<1$ in the STAR experiment [124]). The triangle acceptance can be largely corrected by the mixed-event technique, resulting in two-particle $(\Delta\eta,\Delta\phi)$ correlations, such as the one sketched in Fig. 1, that are dominated by an overall pedestal. The fine structures atop the pedestal are nonflow correlations and a flow modulation along $\Delta\phi$ (with possibly a $\Delta\eta$ dependence). These features have been generally described in the introduction.
The near-side nonflow peaks may be modeled by 2D Gaussians, and the flow modulation is described by a Fourier series. The Fourier coefficients, corresponding to flow harmonics, are typically assumed to be $\Delta\eta$ independent. The two-particle correlations can then be fit with a functional form including Gaussians and Fourier harmonics. This 2D fitting method is data-driven; the nonflow correlation shapes are dictated by the data structure and modeled. The method fits the data in a given centrality or multiplicity bin and does not rely on assumptions using low-multiplicity baseline events. A recent 2D fit study indicates an approximately 40% nonflow in central isobar collisions [122, 123], in line with the central Au+Au data [125] (also see Sect. 3) considering the multiplicity dilution of nonflow by approximately a factor of two.
It is noteworthy that the Fourier flow harmonics can be $\Delta\eta$ dependent, for example, due to flow decorrelations [90, 91, 92, 93]. In addition, the $v_{n}$, although long ranged, can be $\eta$ dependent [87, 73, 88, 89]; such a dependence can result in $\Delta\eta$-dependent Fourier coefficients as well. Including these possible $\eta$-dependencies of the flow harmonics as free parameters in the fit model would not be fruitful as the correlation data would not have sufficient constraining power over these parameters (and those modeling nonflow). Usually the effects due to the possible $\eta$-dependent flow harmonics are assessed as part of the systematic uncertainties [122, 123].
However, the information on the $\Delta\eta$-dependencies of flow harmonics can be obtained from other measurements and can be factored into the 2D fit function. We leave discussions on this to the next section.
@@
The spirit of the one-dimensional (1D) fitting method is similar to that of the 2D fitting method, except that the $\Delta\phi$ dimension is collapsed into a single number $V_{n}(\Delta\eta)$ for each harmonic. The $V_{n}(\Delta\eta)$ is first calculated by Eq. (4) in each $\Delta\eta$ bin. Then $V_{n}(\Delta\eta)$ as a function of $\Delta\eta$ is decomposed into flow and nonflow through the 1D fitting method.
@@
Figure 4 illustrates a typical $V_{2}(\Delta\eta)$ dependence on $\Delta\eta$. It is generally a decreasing function of $\Delta\eta$, primarily because of the short-range nonflow contributions. For tracking detectors, track merging/splitting is sometimes important [120, 105, 122, 123], which creates a dip/peak [^4] HBT/Bose-Einstein correlations and Coulomb interactions can also create dips and peaks at small pair separation [46]. at $\Delta\eta\sim 0$ as indicated by the dashed curve in Fig. 4. When track merging effects are significant, the average $V_{n}$ from full-event analysis can even be smaller than the subevent result which automatically excludes the dip. In such a case, the difference between full-event and subevent $V_{n}$ in the $\Delta\eta$-gap methods would not be a good assessment of systematic uncertainties. On the other hand, when track splitting effects are significant, then the systematic uncertainties would be grossly overestimated.
@@
Figure 4: Illustration of the 1D fit method. The black curve indicates a $V_{2}$ measurement, where the dip at $\Delta\eta\sim 0$ indicated by the dashed curve is caused by track merging in a typical tracking detector. The magenta curve incorporates $\Delta\eta$ convolution from single-particle $v_{2}(\eta)$ and flow decorrelation $1-F_{n}\Delta\eta$. The blue curve indicates the flow component in $V_{2}$ scaled up from the magenta curve to include flow fluctuations, with the scaling factor treated as a fit parameter. The difference between the black curve and the blue curve would be nonflow contribution to the $V_{2}$ measurement. The vertical axis is zero-suppressed and the values are only order-of-magnitude indications.
Compared to the 2D fitting method in Sect. 2.3.1, 1D fitting is more straightforward because the $\Delta\phi$ dimension is collapsed in the $V_{n}$ measurement. It is also easier to include the possible $\Delta\eta$ dependence of flow and flow fluctuations via $\Delta\eta$-dependent $V_{n}$. There are two sources for $\Delta\eta$ dependence of flow as aforementioned:
  • The single particle flow $v_{n}$ can depend on $\eta$ [87, 73, 88, 89], which would result in a $\Delta\eta$-dependent two-particle cumulant $V_{n}(\Delta\eta)$ convoluted from the two single-particle distributions of $v_{2}(\eta_{1})$ and $v_{2}(\eta_{2})$. Such a dependence can be examined, for the case of $n=2$ for example, by four-particle cumulant measurement $v_{2}\{4\}(\eta)$ where nonflow is largely eliminated, and/or by $v_{2}\{{\textsc{zdc}}\}(\eta)$ with respect to the first order harmonic plane of spectator neutrons measured by Zero-Degree Calorimeters (ZDC) which is free of nonflow. Note that $v_{2}\{{\textsc{zdc}}\}$ measures approximately the average $\langle{v_{2}}\rangle$ and $v_{2}^{2}\{4\}\approx\langle{v_{2}}\rangle^{2}-\sigma^{2}$ where $\sigma$ denotes Gaussian width of $v_{2}$ fluctuations, whereas the flow component in the two-particle measurement $V_{2}$ is $v_{2}^{2}\{2\}=\langle{v_{2}}\rangle^{2}+\sigma^{2}$ [29]. As an approximation, one may obtain $v_{2}\{2\}(\eta)$ from $v_{2}\{4\}(\eta)$ and $v_{2}\{{\textsc{zdc}}\}(\eta)$. Alternatively, one may make the reasonable assumption that the effect of $v_{2}$ fluctuations is proportional to $\langle{v_{2}}\rangle$ and apply a scale factor as a free fit parameter to $v_{2}\{4\}(\eta)$ or $v_{2}\{{\textsc{zdc}}\}(\eta)$ to obtain $v_{2}\{2\}(\eta)$.
  • Another source of $\Delta\eta$ dependence is flow decorrelation, $r_{n}(\Delta\eta)\approx 1-F_{n}\Delta\eta$ [90, 91, 92, 93]. Such decorrelations can result from fluctuations of the harmonic flow magnitude as well as harmonic planes over $\Delta\eta$ [126, 127]. The decorrelation parameter $F_{n}$ of a few percent has been measured at RHIC [128] and the LHC [92, 93].
The $\Delta\eta$ dependence of the flow component in two-particle $V_{2}(\Delta\eta)$ is the combination of the two effects above. This is illustrated by the magenta curve in Fig. 4.
The nonflow component in $V_{2}(\Delta\eta)$ can be modeled by typical functional forms like Gaussians in $\Delta\eta$. Fits are applied to $V_{2}(\Delta\eta)$ excluding the track merging $\Delta\eta\sim 0$ region, or alternatively including a negative Gaussian at $\Delta\eta=0$ to model it. The fit function, schematically, is
$V_{2}(\Delta\eta)=\langle{v_{2}}\rangle^{2}\mbox{-shape}\times(1-F_{n}\Delta\eta)\times f+\mbox{nonflow}.$ (15)
The blue curve in Fig. 4 is an illustration of the fitted flow, scaled up from the shape given in magenta. This scaling factor, the $f$ in Eq. (15), is the effect of flow fluctuations, assumed to be independent of $\Delta\eta$ and treated as a fit parameter. The idea of the fit is to have the best description of $V_{2}(\Delta\eta)$ using the known shape of flow and flow fluctuations, and deciding by trial and error on the functional forms for the $\Delta\eta$ dependence of nonflow.